OpenStudy (anonymous):

Solve for x. x2 - 2x - 24 = 0 A. -4, -6 B. -4, 6 C. 2, -6 D. 4, 6

OpenStudy (luigi0210):

What are possible factors of -24?

OpenStudy (jhannybean):

what two numbers add to give 2 and multiply to give -24?

OpenStudy (luigi0210):

*-2

OpenStudy (jhannybean):

-2*:D

OpenStudy (jhannybean):

\[\large x^2-2x-24=0\]\[\large (x-6)(x+4)=0\] solve for x. \[\large (x-6)=0\]\[\large (x+4)=0\]

OpenStudy (luigi0210):

:l

OpenStudy (jhannybean):

...I didn't solve it <_<

OpenStudy (luigi0210):

Here, have a graph

OpenStudy (luigi0210):

it's blue :D

OpenStudy (luigi0210):

You're welcome .-.

OpenStudy (jhannybean):

I was solving a different problem. It's pretty!